Derive a formula for

n

i= i3 a telescoping sum with terms f(I) = i4

1
Expert's answer
2021-03-31T01:58:45-0400

TO DERIVE :

let,

"S=\displaystyle\sum_{i=1}^ni^3" ............(1)


We start off the proof with the "telescoping sum" (or the collapsing sum).


"\displaystyle\sum_{i=1}^n [(i+1)^4-i^4]" SOLVING IN GENERAL FORM


"=(\cancel{2^4}-1^4)+(\cancel{3^4}-\cancel{2^4})+(4^4-\cancel{3^4})\\\\+...........+................+(n+1)^4-\cancel{n^4}\\+...........+................+
\cancel{(n)^4}-\cancel{(n-1)^4}\\
=(n+1)^4-1^4\\
=(n=1)^4-1\\
=((n+1)^2)^2-1\\
=(n^2+2n+1)^2-1"


"\displaystyle\sum_{i=1}^n [(i+1)^4-i^4]" "=n^4+4n^3+6n^2+4n+1" .......................(2)


now,


"\displaystyle\sum_{i=1}^n [(i+1)^4-i^4]" SOLVING IN terms of i



"\u200b\t\n=\\displaystyle\\sum_{i=1}^n[(i^4+4i^3+6i^2+4i+1)-i^4]\\\\\n=\\displaystyle\\sum_{i=1}^n [(4i^3+6i^2+4i+1)] \\\\\n\n=4\\displaystyle\\sum_{i=1}^ni^3+6\\displaystyle\\sum_{i=1}^ni^2+4\\displaystyle\\sum_{i=1}^ni+\\displaystyle\\sum_{i=1}^n1\\\\"

from (1) ..further solving above...


"=4S+6({n(n+1)(2n+1)\over6})+4({n(n+1)\over2})+n"

"=4S+2n^3+n^2+2n^2+n+2n^2+3n\\
=4S+2n^3+5n^2+4n" ................(3)


THUS; equating (2) and (3);


"n^4+4n^3+6n^2+4n=4S+2n^3+5n^2+4n\\
n^4+2n^3+n^2=4S\\
n^2(n^2+2n+1)=4S"


"4S=n^2(n+1)^2\\
S={n^2(n+1)^2\over2^2}"


"\boxed{S=[{n(n+1)\over2}]^2}"

hence drived.